OpenStudy (anonymous):

The equation V=-3000t+2100 can be used to find the value V od a car at the end of t years. e. What is the value of the car at the end of 5 years? f. In how many years is the car worth $12000? g. Is it true that the car will be worth $200 after 7 years? h. How much did the car cost originally?

OpenStudy (anonymous):

e. Plug in... V = -3000(5)+2100 = -$12900. There's something wrong with the equation V=-3000t + 2100... you can't be owing money to your own car... f. The value of the car does not increase, so only if t is negative does the car become 12000, at t= -3.3 years... funny answer yet again. g. V= -3000 (7) + 2100 is NOT 2100. h. 2100 because at t=0, V= 2100